OpenStudy (anonymous):

Which piecewise function coincides with f(x)=|−x−3|2−1?

OpenStudy (anonymous):

f(x)=−(x+3)2−1 when x < 3 f(x)=−(x+3)2−1 when x ≤ -3 f(x)=−(x−3)2−1 when x ≥ 3 f(x)=(x+3)2−1 when x < -3

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

\[|-x-3|^2-1~~=~~\left(|-1||x+3|\right)^2-1~~=~~|x+3|^2-1\] Consider the graph of \(y=(x+3)^2\):  Since every value of \(y\) is positive, you know that \(|x+3|^2=(x+3)^2\). So what this means is \[|-x-3|^2-1=(x+3)^2-1\] If you know the plot of one, you know the plot of the other.

OpenStudy (anonymous):

In terms of a piecewise function, only the last option agrees with the above.